Prove that 

n*P( n -1,n - 1) = P (n, n)


1
Expert's answer
2021-10-25T15:14:45-0400

"P(n,k) = {\frac{n!} {(n-k)!}}"

"P(n,n) = {\frac {n!} {(n-n)!}} = n!"

The point is to prove that "n*P(n-1,n-1) = n!"

"n*P(n-1,n-1) = n*{\frac {(n-1)!} {((n-1)-(n-1))!}}=n*(n-1)!=n!"

The statement has been proven